OpenStudy (anonymous):

Complex numbers: how do i solve the complex equation: z^3 + z1*z = 0, where z1 = sqrt(3) + i?

OpenStudy (anonymous):

z1= 2e^(ipi/6) z(z^2 + z1)=0 z=0 or z = root(-z1) z=0 or z=ie^(ipi/12)

OpenStudy (anonymous):

sorry its 0 or 1.414ie^(ipi/12)

OpenStudy (anonymous):

think there should be three solutions since polynomial has degree 3

OpenStudy (anonymous):

the second soln has a + and -

OpenStudy (anonymous):

\[z(z^2+(\sqrt{3}+i))=o\] \[z=0\] or \[ z=\pm \sqrt{-\sqrt{3}-i}\]

OpenStudy (anonymous):

\[-\sqrt{3}-i=2[cos(\frac{7\pi}{6})+i sin(\frac{7\pi}{6})]=2e^{\frac{7\pi}{6}i}\] unless i made a mistake somewhere.

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

ok so root is \[\sqrt{2}e^{\frac{7\pi}{12}}\]

OpenStudy (anonymous):

or \[\sqrt{2}e^{\frac{19\pi}{12}}\]